OpenStudy (anonymous):

write the slope-intercept form of the equation of the line through the given points> through: (-4,3) and (0,0)... help asap please

OpenStudy (wolfe8):

Do you know how to find slope when given 2 points?

OpenStudy (anonymous):

i know how, i just need the answer, im working on other subjects if that makes sence, but m= y^2-y^1/x^2-x^1

OpenStudy (wolfe8):

Um, I'm not sure why you have powers on there, but it's \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

yeah i know the equation i just need the answer please and thank you :/ ?

OpenStudy (wolfe8):

And then, use the formula \[\left( y _{2}-y _{1} \right)=m\left( x _{2}-x _{1}\right)\]

OpenStudy (anonymous):

AND THE ANSWER IS :)

OpenStudy (wolfe8):

Hm you shouldn't just ask for the answer on this site. It's part of the COC. But I got y=-4x/3-2 1/3

OpenStudy (anonymous):

i know how to do it i have an A in my class, i just dont have time right now cuz im doing other classes :/,

OpenStudy (wolfe8):

Honestly if you got an A it should take you around 2 minutes to do it.

OpenStudy (wolfe8):

I accidentally switched the slope. It's -3/4

OpenStudy (wolfe8):

Bah now I have to re-do it. See it would have been easier if you did it @.@

OpenStudy (anonymous):

yeah i am an A student i just dont got the time for math right now, im writting an essay for my final in english

OpenStudy (anonymous):

:) your so sweet :)

OpenStudy (wolfe8):

I have an essay due tomorrow too. y=-3/4x Dammit I should have read the question better.

OpenStudy (wolfe8):

Sweet?

OpenStudy (anonymous):

(0,_1) and (-4,2)

OpenStudy (anonymous):

and yes your a sweet guy :) \

OpenStudy (wolfe8):

For THAT one you use the formulas I gave you. I don't really feel like doing it...

OpenStudy (anonymous):

:) why are you being lazy uh ^_^, please with a cupcake on top

OpenStudy (anonymous):

cherry on top :)

OpenStudy (wolfe8):

I think it's y=-3x/4-1

OpenStudy (anonymous):

(-2,3) and (1,5)

OpenStudy (wolfe8):

...now you're pushing it D: You gotta work for it yourself too and learn. You're an A student

OpenStudy (anonymous):

:) alright thank for the help whats your real name :)

OpenStudy (wolfe8):

Wolfe. You're welcome

OpenStudy (anonymous):

thank you wolfe...

OpenStudy (wolfe8):

You're welcome Julia